# A Primal-Dual-Assisted Penalty Approach to Bilevel Optimization with Coupled Constraints

### v1

@@ Abstract
Interest in bilevel optimization has grown in recent years, partially due to its applications to tackle challenging machine-learning problems. Several exciting recent works have been centered around developing efficient gradient-based algorithms that can solve bilevel optimization problems with provable guarantees. However, the existing literature mainly focuses on bilevel problems either without constraints, or featuring only simple constraints that do not couple variables across the upper and lower levels, excluding a range of complex applications. Our paper studies this challenging but less explored scenario and develops a (fully) first-order algorithm, which we term BLOCC, to tackle BiLevel Optimization problems with Coupled Constraints. We establish rigorous convergence theory for the proposed algorithm and demonstrate its effectiveness on two well-known real-world applications - hyperparameter selection in support vector machine (SVM) and infrastructure planning in transportation networks using the real data from the city of Seville.
@@ 1 Introduction
where $f:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}$ are respectively the upper-level and lower-level objectives, $g^{c}(x,y):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{c}}$ defines the lower-level CCs, and $\mathcal{X}\subseteq\mathbb{R}^{d_{x}},\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$ are the domain of $x$ and $y$ that are easy to project, such as the Euclidean ball.
The difficulty of solving (1) lies in the coupling between the upper and lower-level problems. To develop efficient methods, previous literature started from the unconstrained BLO problem and solved it by implicit gradient descent (IGD) method [27, 31, 34, 12, 13, 39, 60, 42, 64] and penalty-based methods [44, 43, 60, 40, 41], to name a few. To solve BLO with CCs, AiPOD [68] and GAM [69] have investigated the IGD method under different constraint settings. However, AiPOD only considered equality constraints, and GAM lacked finite-time convergence guarantees. Leveraging a penalty reformulation, [72] developed a Hessian-free method with finite-time convergence. However, a key algorithm step in [72] is a joint projection of the current iterate $(x,y)$ onto the constraint set. This projection, which needs to be carried out at every iteration, can be challenging when the number of variables ($d_{x},d_{y}$) or constraints ($d_{c}$) is large or when $g^{c}(x,y)$ is not jointly convex.
@@
where the penalty term with the constant $\gamma$ controls the distance between $y$ and $y^{*}_{g}(x)$ by penalizing the distance between the lower-level objective $g(x,y)$ and its value function $v(x)$, and the Lagrangian term penalizes the constraint violation of $g^{c}(x,y)$.
@@ 1.1 Main contributions
    In Section 4, we apply the proposed BLOCC algorithm to two real-world applications: SVM model training and planning of a transportation network. By comparison with LV-HBA [72] and GAM [69], we demonstrate the algorithm’s effectiveness.
@@ 1.2 Related works
Methods for BLO without constraints. A cluster of BLO gradient-based approaches gravitates around the implicit gradient descent (IGD) method [54], where the key idea is to approximate the hypergradient by the implicit function theorem. The finite-time convergence of IGD was first established in [27] for unconstrained strongly-convex lower-level problems. Subsequent works [31, 34, 12, 13, 39, 60, 42, 64, 70, 35] improved the convergence rates under various settings. Another cluster of works is based on iterative differentiation (ITD) [48, 23, 52, 59], which estimates the hypergradient by differentiating the entire iterative algorithm used to solve the lower-level problem with respect to the upper-level variables. The finite-time guarantee was first established in [28, 46, 33, 6]. Penalty-based methods have also emerged as a promising approach for BLO. Dated back to [74], this line of works [44, 50, 43, 40, 61, 25, 47] reformulated the original BLO as the single-level problems with various penalty terms and leveraged first-order methods to solve them.
Figure 1: Calculation of $\nabla v(x)$. The blue line is $v(x)$, the red dashed line is calculated by the formulation given in [61, 40], while the yellow line is derived by our BLOCC. It can be seen that $\nabla v(x)$ without the Lagrange multiplier is in the opposite direction of the true gradient.
@@ 2.1 The challenges in BLO with coupled constraints
The first challenge associated with the presence of CCs is that the closed-form expression of the gradient $\nabla v(x)$ without CCs, which is $\nabla_{x}g(x,y_{g}^{*}(x))$ provided in [61, 41, 40, 43], is not applicable. For example, when $g(x,y)=(y-2x)^{2}$ and $g^{c}(x,y)=3x-y$, the optimal lower-level solution is $y_{g}^{*}(x)=3x$ and thus, $v(x)=x^{2}$ with $\nabla v(x)=2x$. However, $\nabla_{x}g(x,y_{g}^{*}(x))=-4x\neq\nabla v(x)$.
The closed form gradient for BLO with CCs should be (8), which considers a Lagrange term that is going to be illustrated later in this paper. In Figure 1, we present the gradient $\nabla v(x)$ without the Lagrange multiplier in [61, 41, 40, 43] and ours with the Lagrange term. In this example, the gradient without the Lagrange term leads to the opposite direction to the true gradient.
The second challenge associated with the presence of CCs is the difficulty of performing the joint projection. If we directly extend the penalty reformulation in [61, 40], we could treat the coupling constraint set $\mathcal{Y}(x)$ as a joint constraint $\{(x,y):g^{c}(x,y)\leq 0\}$, and employ a joint projection of $(x,y)$ to ensure the feasibility. However, this can be computationally inefficient when either the number of variables or constraints is high. Moreover, when $g^{c}(x,y)$ is complex, the projection may not have a closed form and may not even be well-defined if $g^{c}(x,y)$ is not jointly convex.
@@ 2.2 The Lagrangian duality-based penalty reformulation
Domain $\mathcal{X}\subset\mathbb{R}^{d_{x}}$ and $\mathcal{Y}\subset\mathbb{R}^{d_{y}}$ are non-empty, closed, and convex. For any given $x\in\mathcal{X}$, $\mathcal{Y}(x):=\{y\in\mathcal{Y}:g^{c}(x,y)\leq 0\}$ is non-empty.
@@ 2.3 Smoothness of the penalty reformulation
The assumption of the existence of the upper bound for the Lagrange multiplier is mild; see e.g., [32, Theorem 1.6] and [72]. Finding $\nabla v(x)$ is crucial for the design of a gradient-based method to solve $\min_{x\in\mathcal{X}}F_{\gamma}(x)$. Leveraging Lemma 2, the gradient $\nabla F_{\gamma}(x)$ can be obtained by the next lemma.
@@ 3.1 BLOCC algorithm
Suppose that the assumptions in Lemma 3 hold. Run Algorithm 1 with $T_{F}$ and $T_{g}$ sufficiently large such that $(y_{g,t}^{T_{g}},\mu_{g,t}^{T_{g}})$ and $(y_{F,t}^{T_{F}},\mu_{F,t}^{T_{F}})$ are respectively ${\cal O}(\epsilon_{g})$ and ${\cal O}(\epsilon_{F})$-optimal in squared distance as in (11). Set $\eta\leq\frac{1}{l_{F,1}}$ with some $l_{F,1}$ defined in Lemma 3. It then holds that
@@
Using the projected gradient $G_{\eta}=\eta^{-1}(x_{t+1}-x_{t})$ as the convergence metric for constrained optimization problems is standard [26]. The proof of this theorem is available in Appendix D.1. The term ${\cal O}(\gamma^{2}\epsilon_{F}+\gamma^{2}\epsilon_{g})$ arises from estimation errors using the MaxMin Solver. In the next section, we develop a particular MaxMin Solver for Algorithm 1.
@@ 3.2 MaxMin Solver for the BLO with inequality CCs
1:inputs: initial values $\mu_{-1}=\mu_{0}\in\mathbb{R}^{d_{c}}_{+}$, $y_{0}\in\mathcal{Y}$, step sizes $\eta_{1},\eta_{2}>0$, counters $T,T_{y}$. Blue part is the version with acceleration; Red part is without acceleration.
@@
$\mu_{t+\frac{1}{2}}=\mu_{t}+\frac{t-1}{t+2}(\mu_{t}-\mu_{t-1}).$ (17)
@@
Suppose that Assumptions 15 and the conditions in Theorem 2 hold. Let $\gamma>\frac{l_{f,1}}{\alpha_{g}}$, $\epsilon_{g}\leq\frac{C_{\delta_{g}}}{2}\delta_{g}$ and $\epsilon_{F}\leq\frac{C_{\delta_{F}}}{2}\delta_{F}$. If we choose the inner loop $T_{g}={\cal O}(\epsilon_{g}^{-0.5}),T_{F}={\cal O}(\epsilon_{F}^{-0.5})$, and $T_{y}^{g}={\cal O}(\log(\epsilon_{g}^{-1})),T_{y}^{F}={\cal O}(\log(\epsilon_{F}^{-1}))$, then the iterates generated by Algorithm 1 satisfy
$\|\mu_{t,g}^{T_{g}}-\mu^{*}_{g}(x_{t})\|^{2}={\cal O}(\epsilon_{g})\quad\text{and}\quad\|y_{t,g}^{T_{g}}-y^{*}_{g}(x_{t})\|^{2}={\cal O}(\epsilon_{g})$
$\|\mu_{t,F}^{T_{F}}-\mu^{*}_{F}(x_{t})\|^{2}={\cal O}(\epsilon_{F})\quad\text{and}\quad\|y_{t,F}^{T_{F}}-y^{*}_{F}(x_{t})\|^{2}={\cal O}(\epsilon_{F}).$
For solving an $\epsilon$-approximation problem of BLO defined in (4), we solve $\min_{x\in\mathcal{X}}F_{\gamma}(x)$ with $\gamma={\cal O}(\epsilon^{-0.5})$ according to Theorem 1. In this way, to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\|G_{\eta}(x_{t})\|^{2}\leq\epsilon$ where $\frac{1}{T}\sum_{t=0}^{T-1}\|G_{\eta}(x_{t})\|^{2}={\cal O}(\gamma T^{-1}+\gamma^{2}\epsilon_{F}+\gamma^{2}\epsilon_{g})$ by (14), we need $\epsilon_{g},\epsilon_{F}={\cal O}(\epsilon^{2})$, i.e. complexity $\tilde{\cal O}(\epsilon^{-1})$ for the MaxMin solver in Algorithm 2, and $T={\cal O}(\epsilon^{-1.5})$ for the number of iteration in BLOCC (Algorithm 1). Therefore, the overall complexity is $\tilde{\cal O}(\epsilon^{-2.5})$, where $\tilde{\cal O}$ omits the $\log$ terms.
@@
Consider BLO with $\mathcal{Y}=\mathbb{R}^{d_{y}}$ and $g^{c}(x,y)$ defined in (22). Suppose Assumptions 14 and the conditions in Theorem 2 hold. Suppose for any $x\in\mathcal{X}$, there exist constants $s_{\min}$ and $s_{\max}$ such that $0<s_{\min}\leq\sigma_{\min}(g^{c}_{1}(x))\leq\sigma_{\max}(g^{c}_{1}(x))\leq s_{\max}<\infty$. Let $\gamma>\frac{l_{f,1}}{\alpha_{g}}$. If we choose the inner loop $T_{g}={\cal O}(\log(\epsilon_{g}^{-1})),T_{F}={\cal O}(\log(\epsilon_{F}^{-1}))$, and $T_{y}^{g}=T_{y}^{F}=1$ with proper constant stepsizes, then the iterates generated by Algorithm 1 satisfy
$\|\mu_{t,g}^{T_{g}}-\mu^{*}_{g}(x_{t})\|^{2}={\cal O}(\epsilon_{g})\quad\text{and}\quad\|y_{t,g}^{T_{g}}-y^{*}_{g}(x_{t})\|^{2}={\cal O}(\epsilon_{g})$
$\|\mu_{t,F}^{T_{F}}-\mu^{*}_{F}(x_{t})\|^{2}={\cal O}(\epsilon_{F})\quad\text{and}\quad\|y_{t,F}^{T_{F}}-y^{*}_{F}(x_{t})\|^{2}={\cal O}(\epsilon_{F}).$
The proof is available in Appendix D.3. Similar to the previous analysis, we choose $\gamma={\cal O}(\epsilon^{-0.5})$ to solve the equivalent (2a) to $\epsilon$-approximation problem of BLO (4). To achieve $\frac{1}{T}\sum_{t=0}^{T-1}\|G_{\eta}(x_{t})\|^{2}\leq\epsilon$, the number of iteration in BLOCC $T={\cal O}(\epsilon^{-1.5})$ by (14). As the inner MaxMin solver has linear convergence, the overall complexity is $\tilde{\cal O}(\epsilon^{-1.5})$ where $\tilde{\cal O}$ omits the $\log$ terms. We summarized the overall complexity of our BLOCC algorithm in different settings in Table 1.

### v2

@@ Abstract
Interest in bilevel optimization has grown in recent years, partially due to its applications to tackle challenging machine-learning problems. Several exciting recent works have been centered around developing efficient gradient-based algorithms that can solve bilevel optimization problems with provable guarantees. However, the existing literature mainly focuses on bilevel problems either without constraints, or featuring only simple constraints that do not couple variables across the upper and lower-levels, excluding a range of complex applications. Our paper studies this challenging but less explored scenario and develops a (fully) first-order algorithm, which we term BLOCC, to tackle BiLevel Optimization problems with Coupled Constraints. We establish rigorous convergence theory for the proposed algorithm and demonstrate its effectiveness on two well-known real-world applications - hyperparameter selection in support vector machine (SVM) and infrastructure planning in transportation networks using the real data from the city of Seville.
@@ 1 Introduction
where $f:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}$ is the upper-level objective, $g:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}$ is the lower-level objective which is strongly convex, $g^{c}(x,y):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{c}}$ defines the lower-level CCs, and $\mathcal{X}\subseteq\mathbb{R}^{d_{x}},\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$ are the domain of $x$ and $y$ that are easy to project, such as the Euclidean ball.
The difficulty of solving (1) lies in the coupling between the upper and lower-level problems. To develop efficient methods, previous literature started from the unconstrained BLO problem and solved it by implicit gradient descent (IGD) method [28, 32, 35, 12, 13, 41, 62, 44, 67] and penalty-based methods [46, 45, 62, 42, 43], to name a few. To solve BLO with CCs, AiPOD [71] and GAM [72] have investigated the IGD method under different constraint settings. However, AiPOD only considered equality constraints, and GAM lacked finite-time convergence guarantees. Leveraging a penalty reformulation, [75] developed a Hessian-free method with finite-time convergence. However, a key algorithm step in [75] is a joint projection of the current iterate $(x,y)$ onto the coupled constraint set. This projection, required at each iteration, becomes particularly challenging when $g^{c}(x,y)$ is not jointly convex and can be computationally expensive for large-scale problems with a high number of variables ($d_{x},d_{y}$) or constraints ($d_{c}$).
@@
where the penalty constant $\gamma$ controls the distance between $y$ and $y^{*}_{g}(x)$ by penalizing $g(x,y)$ to its value function $v(x)$, and the Lagrangian term penalizes the constraint violation of $g^{c}(x,y)$.
@@ 1.1 Main contributions
    In Section 4, we apply our BLOCC algorithm to two real-world applications: SVM model training and transportation network planning. By comparison with LV-HBA [75] and GAM [72], we demonstrate the algorithm’s effectiveness and its robustness to large-scale problems.
@@ 1.2 Related works
Methods for BLO without constraints. A cluster of BLO gradient-based approaches gravitates around the implicit gradient descent (IGD) method [56], where the key idea is to approximate the hypergradient by the implicit function theorem. The finite-time convergence of IGD was first established in [28] for unconstrained strongly-convex lower-level problems. Subsequent works [32, 35, 12, 13, 41, 62, 44, 67, 73, 36] improved the convergence rates under various settings. Another cluster of works is based on iterative differentiation (ITD) [50, 24, 54, 61], which estimates the hypergradient by differentiating the entire iterative algorithm used to solve the lower-level problem with respect to the upper-level variables. The finite-time guarantee was first established in [29, 48, 34, 6]. Viewing the lower-level problem as a constraint such as in [65], penalty-based methods have also emerged as a promising approach for BLO. Dated back to [77], this line of works [46, 52, 45, 42, 63, 26, 49] reformulated the original BLO as the single-level problems with various penalty terms and leveraged first-order methods to solve them.
@@ 2.1 The challenges in BLO with coupled constraints
The first challenge associated with the presence of CCs is the difficulty to find the descent direction for $x$, which involves finding the closed-form expression of the gradient $\nabla v(x)$. The expression without CCs, $\nabla v(x)=\nabla_{x}g(x,y_{g}^{*}(x))$ provided in [63, 43, 42, 45], is not applicable.
Figure 1: Calculation of $\nabla v(x)$. The blue line is $v(x)$, the the yellow dashed line is calculated by the formulation given in [63, 42], while red dashed line is derived by our BLOCC. It can be seen that $\nabla v(x)$ without the Lagrange multiplier is in the opposite direction of the true gradient.
For example, when $g(x,y)=(y-2x)^{2}$ and $g^{c}(x,y)=3x-y$, the optimal lower-level solution is $y_{g}^{*}(x)=3x$ and thus, $v(x)=x^{2}$ with $\nabla v(x)=2x$. However, $\nabla_{x}g(x,y_{g}^{*}(x))=-4x\neq\nabla v(x)$. The closed form gradient for BLO with CCs should be (8), which considers a Lagrange term that will be illustrated later in this paper. In Figure 1, we present the gradient $\nabla v(x)$ without the Lagrange multiplier in [63, 43, 42, 45] and ours with the Lagrange term. In this example, the gradient without the Lagrange term leads to the opposite direction to the true gradient.
The second challenge associated with the presence of CCs is the difficulty of performing the joint projection. If we directly extend the penalty reformulation in [63, 42], we could treat the coupling constraint set $\mathcal{Y}(x)$ as a joint constraint $\{(x,y):g^{c}(x,y)\leq 0\}$, and employ a joint projection of $(x,y)$ to ensure the feasibility. However, this can be computationally inefficient when the problem is large-scale, i.e. the number of variables or constraints is large. The detailed analysis of computational cost can be seen in Appendix H. Moreover, when $g^{c}(x,y)$ is complex, the projection may not have a closed form and may not even be well-defined if $g^{c}(x,y)$ is not jointly convex.
@@ 2.2 The Lagrangian duality-based penalty reformulation
Domain $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ and $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$ are non-empty, closed, and convex. For any given $x\in\mathcal{X}$, $\mathcal{Y}(x):=\{y\in\mathcal{Y}:g^{c}(x,y)\leq 0\}$ is non-empty.
@@ 2.3 Smoothness of the penalty reformulation
The assumption of the existence of the upper bound for the Lagrange multiplier is a consequence of the LICQ condition [70, Theorem 1]. This assumption is mild and traditional [75]. Finding $\nabla v(x)$ is crucial for the design of a gradient-based method to solve $\min_{x\in\mathcal{X}}F_{\gamma}(x)$. Leveraging Lemma 2, the gradient $\nabla F_{\gamma}(x)$ can be obtained by the next lemma.
@@ 3.1 BLOCC algorithm
Suppose that the assumptions in Lemma 3 hold. Run Algorithm 1 with some effective inner MaxMin solver to find $(y_{g,t}^{T_{g}},\mu_{g,t}^{T_{g}})$ and $(y_{F,t}^{T_{F}},\mu_{F,t}^{T_{F}})$ respectively ${\cal O}(\epsilon_{g})$ and ${\cal O}(\epsilon_{F})$-optimal in squared distance as in (11). Set $\eta\leq\frac{1}{l_{F,1}}$ with some $l_{F,1}$ defined in Lemma 3. It then holds that
@@
The proof is available in Appendix D.1. Using the projected gradient $G_{\eta}=\eta^{-1}(x_{t+1}-x_{t})$ as the convergence metric for constrained optimization problems is standard [27]. The term ${\cal O}(\gamma^{2}\epsilon_{F}+\gamma^{2}\epsilon_{g})$ arises from estimation errors using the MaxMin Solver. Although the inner oracle can be any one that achieves (11), we value the computational effectiveness and therefore present a particular solver Algorithm 2 in the following section.
@@ 3.2 MaxMin Solver for the BLO with inequality CCs
1:inputs: initial points $y_{0}$, $\mu_{0}$; stepsizes $\eta_{1},\eta_{2}$; counters $T,T_{y}$. Blue part is the version with acceleration; and red part is without.
@@
$\mu_{t+\frac{1}{2}}=\mu_{t}+\frac{t-1}{t+2}(\mu_{t}-\mu_{t-1}),\quad\text{with}~{}\mu_{-1}=\mu_{0}.$ (17)
@@
Suppose that Assumptions 15 and the conditions in Theorem 2 hold. Let $\gamma>\frac{l_{f,1}}{\alpha_{g}}$, $\epsilon_{g}\leq\frac{C_{\delta_{g}}}{2}\delta_{g}$ and $\epsilon_{F}\leq\frac{C_{\delta_{F}}}{2}\delta_{F}$. If we choose Algorithm 2 with acceleration as the inner loop and input $T_{g}={\cal O}(\epsilon_{g}^{-0.5}),T_{F}={\cal O}(\epsilon_{F}^{-0.5})$, and $T_{y}^{g}={\cal O}(\ln(\epsilon_{g}^{-1})),T_{y}^{F}={\cal O}(\ln(\epsilon_{F}^{-1}))$ with proper constant stepsizes in Remark 2, then the iterates generated by the Algorithm 1 satisfy (11):
$\|(y_{g,t}^{T_{g}},\mu_{g,t}^{T_{g}})-(y_{g}^{*}(x_{t}),\mu_{g}^{*}(x_{t}))\|^{2}={\cal O}(\epsilon_{g});~{}~{}~{}\|(y_{F,t}^{T_{F}},\mu_{F,t}^{T_{F}})-(y_{F}^{*}(x_{t}),\mu_{F}^{*}(x_{t}))\|^{2}={\cal O}(\epsilon_{F}).$
For solving an $\epsilon$-approximation problem of BLO defined in (4), we solve $\min_{x\in\mathcal{X}}F_{\gamma}(x)$ with $\gamma={\cal O}(\epsilon^{-0.5})$ according to Theorem 1. In this way, to achieve $\frac{1}{T}\sum_{t=0}^{T-1}\|G_{\eta}(x_{t})\|^{2}={\cal O}(\gamma T^{-1}+\gamma^{2}\epsilon_{F}+\gamma^{2}\epsilon_{g})\leq\epsilon$ by (14), we need $\epsilon_{g},\epsilon_{F}={\cal O}(\epsilon^{2})$, i.e. complexity $\tilde{\cal O}(\epsilon^{-1})$ for the MaxMin solver in Algorithm 2, and $T={\cal O}(\gamma\epsilon^{-1})={\cal O}(\epsilon^{-1.5})$ for the number of iteration in BLOCC (Algorithm 1). Therefore, the overall complexity is $\tilde{\cal O}(\epsilon^{-2.5})$, where $\tilde{\cal O}$ omits the $\ln$ terms.
@@
Consider BLO with $\mathcal{Y}=\mathbb{R}^{d_{y}}$ and $g^{c}(x,y)$ defined in (22). Suppose Assumptions 14 and the conditions in Theorem 2 hold. Suppose for any $x\in\mathcal{X}$, there exist constants $s_{\min}$ and $s_{\max}$ such that $0<s_{\min}\leq\sigma_{\min}(g^{c}_{1}(x))\leq\sigma_{\max}(g^{c}_{1}(x))\leq s_{\max}<\infty$. Let $\gamma>\frac{l_{f,1}}{\alpha_{g}}$. If we choose Algorithm 2 without acceleration as the inner loop and input $T_{g}={\cal O}(\ln(\epsilon_{g}^{-1})),T_{F}={\cal O}(\ln(\epsilon_{F}^{-1}))$, and $T_{y}^{g}=T_{y}^{F}=1$ with proper constant stepsizes in Remark 3, then the iterates generated by Algorithm 1 satisfy (11):
$\|(y_{g,t}^{T_{g}},\mu_{g,t}^{T_{g}})-(y_{g}^{*}(x_{t}),\mu_{g}^{*}(x_{t}))\|^{2}={\cal O}(\epsilon_{g});~{}~{}~{}\|(y_{F,t}^{T_{F}},\mu_{F,t}^{T_{F}})-(y_{F}^{*}(x_{t}),\mu_{F}^{*}(x_{t}))\|^{2}={\cal O}(\epsilon_{F}).$
The proof is available in Appendix D.3. Similar to the previous analysis, we choose $\gamma={\cal O}(\epsilon^{-0.5})$ to solve the equivalent (2a) to $\epsilon$-approximation problem of BLO (4). To achieve $\frac{1}{T}\sum_{t=0}^{T-1}\|G_{\eta}(x_{t})\|^{2}\leq\epsilon$, the number of iteration in BLOCC $T={\cal O}(\gamma\epsilon^{-1})={\cal O}(\epsilon^{-1.5})$ by (14). As the inner MaxMin solver convergences linearly, the overall complexity is $\tilde{\cal O}(\epsilon^{-1.5})$ where $\tilde{\cal O}$ omits the $\ln$ terms. We summarized the overall iteration complexity of our BLOCC algorithm in different settings in Table 1.